1.0345 g of palmitic acid dissolved in 8.7545 g of stearic acid. Calculate the freezing point of the solution. The molar mass of palmitic acid is 256.48 g/ mol. The molal freezing point constant , Kf for stearic acid is 4.5 degrees C/ m and pure stearic acid freezes at 69.3 degrees Celsius.
Answer #1

Moles of palmitic acid = mass / molar mass

= 1.0345 / 256.48

= 4.033*10^-3

Molality = moles / mass of solvent (in kg)

= 4.033*10^-3 / ( 8.7545*10^-3)

= 0.461

∆T = kf*m

= 4.5 * 0.461

= 2.0745

T = T° - ∆T

= 69.3 - 2.0745

= 67.23 °C

If you have any query please comment

If you satisfied with the solution please like it thankxx